OpenStudy (anonymous):

If f(x) = 5x + 4 and g(x) = 6x – 3, what is f(g(x))?

OpenStudy (anonymous):

is easy to find

OpenStudy (anonymous):

you already got the value for g(x) and f(x)

OpenStudy (anonymous):

first plug the value of f(x) which is f(x)=5x+4 then second put the value of g(x) inside the f(x), f(g(x))= 5(6x-3)+4 and i leave the rest for you to solve it

OpenStudy (anonymous):

so its f(g(x)) = 5(6x-3)+4 f(g(x)) = 30x-15 + 4 f(g(x)) = 30x-11

OpenStudy (anonymous):

that pretty much a pre-calculus problem which is easy to solve if u are already in calculus